How many ways can you arrange 5 employees in a circular table with only three seats?

Answer by sudhanshu_kmr(1152)


no. of ways to choose 3 employee out of 5 = 5C3 = 10
no. of ways to arrange 3 employee in a circular table = (3-1)! =2

total no. of ways = 10 *2 = 20